If a rock falls from a height of 90 meters above the ground, the height H (in meters) after x seconds can be approximated using the formula H = 90 - 4.9x2. What is the height of the rock after 4 seconds?
1
Expert's answer
2020-09-06T12:49:32-0400

From the formula given above;

H = 90 - 4.9x^2,

The height of the rock can be obtained just by replacing the given value of x = 4 seconds into the formula above i.e.

H = 90 - 4.9 (4)^2

H = 90 - (4.9 × 16)

H = 90 - 78.4

H = 11.6

Therefore the height of the rock after 4 seconds is 11.6 meters.